# declare library=forked edges

### v1

@@ 3.2 WM IP Checkers
After watermarking the LLM outputs, the clouds can use IP checkers to determine whether an adversary model has utilized watermarked outputs to replicate the behavior of the cloud-hosted model. This is typically done by checking whether the generated text from the model is watermarked. IP checkers are generally divided into two main approaches. First, some IP checkers categorize each token as green or red and then apply statistical tests, such as z-scores, p-values, or Jensen-Shannon divergence, to assess whether the proportion of green tokens exceeds a specified threshold. If this threshold is met, the text is flagged as watermarked [9, 50, 59]. Second, other IP checkers assess the alignment between generated tokens and pseudo-random sequence, checking whether each token matches the corresponding value in the sequence. Depending on a WM technique, the cloud will apply its specific detection function to assess the output. If the statistical tests or alignment scores exceed certain thresholds, the output is considered watermarked. If the proportion of detected watermarks is sufficiently high, the model may be flagged as potentially stolen.
@@ 3.3 Characteristics of an Effective WM
A variety of WMs in LLMs have been introduced, while effective WMs in LLMs share key characteristics: First, WMs should minimally impact text quality, ensuring they are not noticeable to users. Quality can be assessed using metrics like perplexity, BLEU score, or accuracy on tasks like MMLU or sentiment classification [92]. Second, WMs must be easily detectable by IP checkers, enabling reliable origin tracing with high detection rates. Third, WMs should be robust to WM removal attacks that attempt to remove watermarked patterns to bypass detection. By satisfying these characteristics, WMs in LLMs can effectively balance security, model utility, and usability in real-world deployments.
@@ 3.4 Attacks against WMs
WM Removal Attacks [26, 93, 94, 27]. The attacker’s goal is to remove embedded signatures and patterns from a watermarked text using paraphrasing techniques such as synonym replacement, sentence reordering, or sentence matching to minimize detectable traces of the WM. These attacks aim to generate semantically equivalent but syntactically modified text to evade WM detection mechanisms. In addition, some attacks [26] propose leveraging a high-quality oracle model to preserve output quality while removing the WM, making detection even more challenging. However, these approaches often require additional training to effectively rephrase sentences or entire paragraphs, which can introduce subtle shifts in semantic meaning, degrade text quality, and significantly increase computational complexity.
Spoofing Attacks [27, 48, 95, 96, 84]. The attacker aims to produce text that is falsely detected as watermarked, thereby nullifying the value of WMs [27]. These attacks undermine the integrity of the WM by making it difficult to distinguish between legitimate watermarked content and counterfeited outputs. These approaches may harm the reputation of model providers, for example, if inappropriate or harmful text contents are falsely attributed to them. However, these attacks work only with certain types of WMs, and to be successful, they typically require a large number of queries to collect sufficient watermarked text for estimating the WM’s distribution and reconstructing the WM rules, associating with computational overheads.
Figure 4: US LLM Market Trend from 2020 to 2030 [97].
@@
ChatGPT and GPT family [2] OpenAI Salesforce integrated its AI with OpenAI for CRM tools [98]. Duolingo introduced Duolingo Max powered by GPT-4 for better learning experience [99]. Instacart added ChatGPT to its Ask Instacart chatbot for better customer interaction [100]. Expedia powered its chatbots by ChatGPT for comprehensive services [101]. Morgan Stanley introduced AskResearchGPT for staff support [102]. Stripe uses GPT-4 to improve user experience and fight fraud [103].
@@
Microsoft Azure Language [32] Microsoft Azure AT&T leveraged Azure OpenAI to automate processes and improve customer experiences [32]. Vodafone improved services with TOBi AI for personalized, multi-channel support [104]. Volvo Group used AI for document translation to streamline invoice & claims processing[104]. TotalEnergies launched Microsoft Copilot for an AI chat solution using internal data [105]. Providence created ProvARIA based on Azure OpenAI to direct patient care messages [105]. Grupo Bimbo developed a copilot for employee queries on policies and risk management [105].
@@
LLaMA family [33] Meta Zoom utilized LLaMA-2 to create its AI Companion for meeting summaries, next-step highlights, and presentation tips [106]. Mathpresso built MathGPT with LLaMA for personalized math learning [106]. Goldman Sachs used LLaMA to improve service, document review, code generation[106]. Nomura Holdings used LLaMA to enhance customer service and document analysis[106]. EPFL developed Meditron with LLaMA-2 to aid clinical decision-making and diagnosis[106].
@@
Google Cloud AI-Language (Gemini, Bard, PaLM, etc.) [107] Google Lilt integrated Google Translate into its platform for advanced translation capabilities [108]. Wipro with Google Cloud created the Gemini Experience Zone for AI experimentation [109]. General Motors deployed a virtual assistant using Google Cloud’s conversational AI for better intent recognition [110]. Gojek launched ”Dira by GoTo AI”, an voice assistant for GoPay service[110]. GroupBy built a Search and Discovery Platform using Vertex AI to boost retail sales[110]. Motorola AI leveraged Gemini for conversation summaries and search[110].
@@
Amazon Comprehend [111] Amazon AWS Vibes used the platform for phrase extraction, sentiment analysis, and topic modeling [112]. Gallup applied Targeted Sentiment to improve survey response reporting[112]. Zillow built speech analytics to enhance customer and service support[112]. Siemens leveraged AWS LLMs to automate its surveying program[112].
@@
IBM Watson NLU [113] IBM Watson SherloQ used IBM Watson NLU to analyze search behavior and optimize ads [114]. University of Auckland deployed IBM Watson Assistant based on a conversational AI solution for handling repetitive questions [114]. Mushi Lab created Clearscope to optimize content with insights & recommendations [114]. Global-Regulation built a global law search engine for easier legal information access[114].
@@
Claude [3] Anthropic Notion integrated Claude for content generation, summarization, and workflow automation [115]. Decagon utilized Claude for personalized, 24/7 customer service [115]. GitLab employed Claude to enhance coding assistance & software development[115]. Headstart leveraged Claude to accelerate their software development process[115].
@@ 4.1 WMs in LLM Productions and Challenges
Given the increasing trend in the LLM market, along with potential risks associated with their deployment in real-world scenarios, as discussed in Section 2, it is essential to utilize WMs to protect IP of LLM outputs. However, the adoption of WMs in LLM productions remains limited. For instance, OpenAI has developed a WM system for ChatGPT but has opted not to implement it due to concerns over ambiguous potential penalties for users and feedback suggesting that service usage might decline if WMs were applied [116]. The uncertainty from providers and users around adopting WMs in LLMs stems from a lack of understanding of WM effectiveness and how different LLMs interact with IP checkers for a given WM.
@@ 4.4.1 Baselines
We carry out the validation on a wide range of WMs, LLMs, and IP classifiers. For LLMs, we use four state-of-the-art (SOTA) pre-trained LLMs, including LLaMA-2 with 7B parameters [117], Mistral with 7B parameters[118], OPT with 6.7B parameters[119] and Falcon with 7B parameters[120]. For WMs, we use seven SOTA WM schemes, including 1) KGW [9], 2) EXP [25], 3) SIR [50], 4) SemStamp [77], 5) Unigram [59], 6) Adaptive [53], and 7) UPV [23]. These WMs are representative WM schemes outlined in Section 3, with further details provided in Table VII (Appendix).
For IP classifiers, we explore a range of options, including 1) a non-transformer LSTM model [121]; 2) an encoder-based transformer model ALBERT [122], with 11.8M parameters; 3) another larger encoder-based transformer DeBERTa [123] with 86M parameters; 4) a decoder-based transformer GPT-2 [31] with 125M parameters; 5) a decoder-only model Galactica [124] with 125M parameters; 6) a small version of OPT [119] with 125M parameters; 7) the recently developed SmolLM [125] with 135M parameters; and 8) an encoder-decoder model T5 [40], with 60.5M parameters. These models were selected based on the following criteria: 1) inclusion of non-transformer and transformer-based architectures, 2) within transformer-based models, coverage of all major categories outlined in Section 2, and 3) consideration of model complexity as determined by parameter size.
To further assess the effectiveness of WMs, we test them against SOTA attacks, including 1) a WM removal attack [26], referred to as WMremoval, which removes WMs by perturbing quality-preserving WM outputs to produce high-quality non-WMed outputs; 2) Dipper [93], which paraphrases paragraphs through content reordering and lexical changes without appreciably modifying the text semantics, and 3) Substitution attack [94], which randomly substitutes target words with their synonyms, considering the surrounding context and using WordNet [126]. We adopt the default settings reported in these papers.
@@ 4.4.2 Datasets and Evaluation Tasks
We randomly select $10,000$ training samples and $2,000$ test samples from the benchmark C4 dataset [127]. Each sample, a news-like string, is truncated to $200$ tokens as a prompt, with the next $200$ tokens used as the “baseline” completion. Tasks include text generation as the primary task and massive multi-task language understanding (MMLU) [92] as a downstream task, assessing the impact of WMs on LLM utility. MMLU has $57$ subjects in STEM, humanities, and social sciences, covering various difficulty levels to evaluate world knowledge and problem-solving skills.
@@ 4.4.3 Evaluation Metrics
where $precision$ is defined as the number of true positive results divided by the total number of all samples predicted by the model, including incorrect predictions, and $recall$ is the number of true positive results divided by the total number of all actual instances. $WM$ and $NW$ indicates the score from WM and no WM settings, respectively.
In addition, to evaluate the model utility comprehensively, we consider 1) Perplexity for the text generation task, measuring how well a model predicts next tokens [128] and 2) Average accuracy across topics for the MMLU downstream task. To further assess text quality, we provide qualitative examples comparing a prompt, the original output without WMs, and corresponding outputs with different WMs to offer insights into WM effectiveness.
@@ 5.2 Robustness to Removal Attacks
WMs, regardless of their category, generally face challenges related to robustness. Given the complex manipulation behaviors of users and adversaries in the real-world, it is difficult to create a WM that can resist to modifications such as paraphrasing, lexical changes, or reordering. As also claimed in [26], current strong WMs fail to offer the desired level of security, especially against sophisticated attacks.
@@ 5.6 Explore WMs and Alignment in LLMs
Alignment refers to ensuring that LLMs generate outputs consistent with standard ethical and operational guidelines [129]. However, WMs can interfere with this alignment by altering the generation or model training process, potentially reducing the model’s adherence to its pre-trained alignment. This issue becomes particularly critical in sensitive domains where the model’s behaviors must align with strict safety or ethical standards. The impact of WMs on LLM alignment demands careful investigation, as disruptions in logits, sampling, or model weights could cause semantic drift or unintended biases in LLM outputs.

### v2

@@ 3.2 WM IP Checkers
After watermarking the LLM outputs, the clouds can use IP checkers to determine whether an adversary model has utilized watermarked outputs to replicate the behavior of the cloud-hosted model. This is typically done by checking whether the generated text from the model is watermarked. IP checkers are generally divided into two main approaches. First, some IP checkers categorize each token as green or red and then apply statistical tests, such as z-scores, p-values, or Jensen-Shannon divergence, to assess whether the proportion of green tokens exceeds a specified threshold. If this threshold is met, the text is flagged as watermarked [kirchenbauer2023watermark, liu2024semanticinvariantrobustwatermark, zhao2024provable]. Second, other IP checkers assess the alignment between generated tokens and pseudo-random sequence, checking whether each token matches the corresponding value in the sequence. Depending on a WM technique, the cloud will apply its specific detection function to assess the output. If the statistical tests or alignment scores exceed certain thresholds, the output is considered watermarked. If the proportion of detected watermarks is sufficiently high, the model may be flagged as potentially stolen.
@@ 3.3 Characteristics of an Effective WM
A variety of WMs in LLMs have been introduced, while effective WMs in LLMs share key characteristics: First, WMs should minimally impact text quality, ensuring they are not noticeable to users. Quality can be assessed using metrics like perplexity, BLEU score, or accuracy on tasks like MMLU or sentiment classification [hendrycks2021measuringmassivemultitasklanguage]. Second, WMs must be easily detectable by IP checkers, enabling reliable origin tracing with high detection rates. Third, WMs should be robust to WM removal attacks that attempt to remove watermarked patterns to bypass detection. By satisfying these characteristics, WMs in LLMs can effectively balance security, model utility, and usability in real-world deployments.
@@ 3.4 Attacks against WMs
WM Removal Attacks [zhang2023watermarks, krishna2024paraphrasing, pan2024markllm, pang2024attacking]. The attacker’s goal is to remove embedded signatures and patterns from a watermarked text using paraphrasing techniques such as synonym replacement, sentence reordering, or sentence matching to minimize detectable traces of the WM. These attacks aim to generate semantically equivalent but syntactically modified text to evade WM detection mechanisms. In addition, some attacks [zhang2023watermarks] propose leveraging a high-quality oracle model to preserve output quality while removing the WM, making detection even more challenging. However, these approaches often require additional training to effectively rephrase sentences or entire paragraphs, which can introduce subtle shifts in semantic meaning, degrade text quality, and significantly increase computational complexity.
Spoofing Attacks [pang2024attacking, jovanovic2024watermark, zhang2024large, sadasivan2023can, gu2023learnability]. The attacker aims to produce text that is falsely detected as watermarked, thereby nullifying the value of WMs [pang2024attacking]. These attacks undermine the integrity of the WM by making it difficult to distinguish between legitimate watermarked content and counterfeited outputs. These approaches may harm the reputation of model providers, for example, if inappropriate or harmful text contents are falsely attributed to them. However, these attacks work only with certain types of WMs, and to be successful, they typically require a large number of queries to collect sufficient watermarked text for estimating the WM’s distribution and reconstructing the WM rules, associating with computational overheads.
Figure 4: US LLM Market Trend from 2020 to 2030 [industry].
@@
ChatGPT and GPT family [OpenAI] OpenAI Salesforce integrated its AI with OpenAI for CRM tools [salesforce2023einstein]. Duolingo introduced Duolingo Max powered by GPT-4 for better learning experience [duolingo2023max]. Instacart added ChatGPT to its Ask Instacart chatbot for better customer interaction [wsj2023instacart]. Expedia powered its chatbots by ChatGPT for comprehensive services [expedia2023chatgpt]. Morgan Stanley introduced AskResearchGPT for staff support [morganstanley2023openai]. Stripe uses GPT-4 to improve user experience and fight fraud [stripe2023openai].
@@
Microsoft Azure Language [MicrosoftAzureAI] Microsoft Azure AT&T leveraged Azure OpenAI to automate processes and improve customer experiences [MicrosoftAzureAI]. Vodafone improved services with TOBi AI for personalized, multi-channel support [azure2023openai]. Volvo Group used AI for document translation to streamline invoice & claims processing[azure2023openai]. TotalEnergies launched Microsoft Copilot for an AI chat solution using internal data [microsoft2024ai]. Providence created ProvARIA based on Azure OpenAI to direct patient care messages [microsoft2024ai]. Grupo Bimbo developed a copilot for employee queries on policies and risk management [microsoft2024ai].
@@
LLaMA family [HuggingFaceMetaLLaMA] Meta Zoom utilized LLaMA-2 to create its AI Companion for meeting summaries, next-step highlights, and presentation tips [meta2024llama]. Mathpresso built MathGPT with LLaMA for personalized math learning [meta2024llama]. Goldman Sachs used LLaMA to improve service, document review, code generation[meta2024llama]. Nomura Holdings used LLaMA to enhance customer service and document analysis[meta2024llama]. EPFL developed Meditron with LLaMA-2 to aid clinical decision-making and diagnosis[meta2024llama].
@@
Google Cloud AI-Language (Gemini, Bard, PaLM, etc.) [GoogleCloudNLP] Google Lilt integrated Google Translate into its platform for advanced translation capabilities [lilt2024googletranslate]. Wipro with Google Cloud created the Gemini Experience Zone for AI experimentation [wipro2024news]. General Motors deployed a virtual assistant using Google Cloud’s conversational AI for better intent recognition [googlecloud2024aiusecases]. Gojek launched ”Dira by GoTo AI”, an voice assistant for GoPay service[googlecloud2024aiusecases]. GroupBy built a Search and Discovery Platform using Vertex AI to boost retail sales[googlecloud2024aiusecases]. Motorola AI leveraged Gemini for conversation summaries and search[googlecloud2024aiusecases].
@@
Amazon Comprehend [satyanarayana2020sentimental] Amazon AWS Vibes used the platform for phrase extraction, sentiment analysis, and topic modeling [aws2024comprehend]. Gallup applied Targeted Sentiment to improve survey response reporting[aws2024comprehend]. Zillow built speech analytics to enhance customer and service support[aws2024comprehend]. Siemens leveraged AWS LLMs to automate its surveying program[aws2024comprehend].
@@
IBM Watson NLU [IBMNLP] IBM Watson SherloQ used IBM Watson NLU to analyze search behavior and optimize ads [ibm2024casestudies]. University of Auckland deployed IBM Watson Assistant based on a conversational AI solution for handling repetitive questions [ibm2024casestudies]. Mushi Lab created Clearscope to optimize content with insights & recommendations [ibm2024casestudies]. Global-Regulation built a global law search engine for easier legal information access[ibm2024casestudies].
@@
Claude [AnthropicClaude] Anthropic Notion integrated Claude for content generation, summarization, and workflow automation [anthropic2024customers]. Decagon utilized Claude for personalized, 24/7 customer service [anthropic2024customers]. GitLab employed Claude to enhance coding assistance & software development[anthropic2024customers]. Headstart leveraged Claude to accelerate their software development process[anthropic2024customers].
@@ 4.1 WMs in LLM Productions and Challenges
Given the increasing trend in the LLM market, along with potential risks associated with their deployment in real-world scenarios, as discussed in Section 2, it is essential to utilize WMs to protect IP of LLM outputs. However, the adoption of WMs in LLM productions remains limited. For instance, OpenAI has developed a WM system for ChatGPT but has opted not to implement it due to concerns over ambiguous potential penalties for users and feedback suggesting that service usage might decline if WMs were applied [chatgptWM]. The uncertainty from providers and users around adopting WMs in LLMs stems from a lack of understanding of WM effectiveness and how different LLMs interact with IP checkers for a given WM.
@@ 4.4.1 Baselines
We carry out the validation on a wide range of WMs, LLMs, and IP classifiers. For LLMs, we use four state-of-the-art (SOTA) pre-trained LLMs, including LLaMA-2 with 7B parameters [touvron2023llama2openfoundation], Mistral with 7B parameters[jiang2023mistral7b], OPT with 6.7B parameters[zhang2022opt] and Falcon with 7B parameters[almazrouei2023falcon]. For WMs, we use seven SOTA WM schemes, including 1) KGW [kirchenbauer2023watermark], 2) EXP [kuditipudi2023robust], 3) SIR [liu2024semanticinvariantrobustwatermark], 4) SemStamp [hou2023semstamp], 5) Unigram [zhao2024provable], 6) Adaptive [liu2024adaptive], and 7) UPV [liu2023unforgeable]. These WMs are representative WM schemes outlined in Section 3, with further details provided in Table VII (Appendix).
For IP classifiers, we explore a range of options, including 1) a non-transformer LSTM model [hochreiter1997long]; 2) an encoder-based transformer model ALBERT [DBLP:journals/corr/abs-1909-11942], with 11.8M parameters; 3) another larger encoder-based transformer DeBERTa [he2021deberta] with 86M parameters; 4) a decoder-based transformer GPT-2 [radford2019language] with 125M parameters; 5) a decoder-only model Galactica [taylor2022galactica] with 125M parameters; 6) a small version of OPT [zhang2022opt] with 125M parameters; 7) the recently developed SmolLM [huggingface2024smollm] with 135M parameters; and 8) an encoder-decoder model T5 [raffel2020exploring], with 60.5M parameters. These models were selected based on the following criteria: 1) inclusion of non-transformer and transformer-based architectures, 2) within transformer-based models, coverage of all major categories outlined in Section 2, and 3) consideration of model complexity as determined by parameter size.
To further assess the effectiveness of WMs, we test them against SOTA attacks, including 1) a WM removal attack [zhang2023watermarks], referred to as WMremoval, which removes WMs by perturbing quality-preserving WM outputs to produce high-quality non-WMed outputs; 2) Dipper [krishna2024paraphrasing], which paraphrases paragraphs through content reordering and lexical changes without appreciably modifying the text semantics, and 3) Substitution attack [pan2024markllm], which randomly substitutes target words with their synonyms, considering the surrounding context and using WordNet [miller1995wordnet]. We adopt the default settings reported in these papers.
@@ 4.4.2 Datasets and Evaluation Tasks
We randomly select $10,000$ training samples and $2,000$ test samples from the benchmark C4 dataset [dodge2021documentinglargewebtextcorpora]. Each sample, a news-like string, is truncated to $200$ tokens as a prompt, with the next $200$ tokens used as the “baseline” completion. Tasks include text generation as the primary task and massive multi-task language understanding (MMLU) [hendrycks2021measuringmassivemultitasklanguage] as a downstream task, assessing the impact of WMs on LLM utility. MMLU has $57$ subjects in STEM, humanities, and social sciences, covering various difficulty levels to evaluate world knowledge and problem-solving skills.
@@ 4.4.3 Evaluation Metrics
where precisionprecision is defined as the number of true positive results divided by the total number of all samples predicted by the model, including incorrect predictions, and recallrecall is the number of true positive results divided by the total number of all actual instances. $WM$ and $NW$ indicates the score from WM and no WM settings, respectively.
In addition, to evaluate the model utility comprehensively, we consider 1) Perplexity for the text generation task, measuring how well a model predicts next tokens [mikolov2011empirical] and 2) Average accuracy across topics for the MMLU downstream task. To further assess text quality, we provide qualitative examples comparing a prompt, the original output without WMs, and corresponding outputs with different WMs to offer insights into WM effectiveness.
@@ 5.2 Robustness to Removal Attacks
WMs, regardless of their category, generally face challenges related to robustness. Given the complex manipulation behaviors of users and adversaries in the real-world, it is difficult to create a WM that can resist to modifications such as paraphrasing, lexical changes, or reordering. As also claimed in [zhang2023watermarks], current strong WMs fail to offer the desired level of security, especially against sophisticated attacks.
@@ 5.6 Explore WMs and Alignment in LLMs
Alignment refers to ensuring that LLMs generate outputs consistent with standard ethical and operational guidelines [shen2023large]. However, WMs can interfere with this alignment by altering the generation or model training process, potentially reducing the model’s adherence to its pre-trained alignment. This issue becomes particularly critical in sensitive domains where the model’s behaviors must align with strict safety or ethical standards. The impact of WMs on LLM alignment demands careful investigation, as disruptions in logits, sampling, or model weights could cause semantic drift or unintended biases in LLM outputs [verma2025watermarking].
@@ 6 Conclusion
Impacts of other WMs on IP Classifiers. As we applied multiple WMs to reassure our assessment, the cross-model IP classifier results of Unigram, Adaptive and UPV WMs are shown in Fig. 12. Similar insights could be drawn from the figure, reaffirming our observation of the study.
(a) Unigram
(b) Adaptive
(c) UPV
Figure 12: $F1$ scores as a function of different IP classifiers for different WMs, compared with No WM settings (Additional to results in Figure 7).
APIs for Deployment of Major LLM Providers. To provide context on the current deployment status of LLMs, we present Table VI, which compares various real-world APIs, outlining their applications, advantages, and limitations. Leading providers like OpenAI, Microsoft, Meta, Hugging Face, Google, and Amazon offer a range of LLM capabilities, including NLP, sentiment analysis, code generation, and model fine-tuning. While each API offers strengths such as scalability, integration, and community support, they also come with challenges like high resource demands, usage-based pricing, and limited customization options.
TABLE VI: LLM Real-world APIs.
API Name Provider Access Type Applications Advantages Constraints
ChatGPT and GPT family [OpenAI] OpenAI API NLP, text generation, questions and answers, chatbots State-of-the-art LLMs, versatility, efficiency, personalization, cost-effective Lack of real-Time knowledge, cost for high usage
Microsoft Azure Language [MicrosoftAzureAI] Microsoft Azure API Sentiment analysis, entity recognition, language understanding Comprehensive APIs, scalability, integration with Azure ecosystem Pricing based on usage, limited offline capabilities
LLaMA family [HuggingFaceMetaLLaMA] Meta Open source Text/code generation, text summarization, translation, recommendations, chatbots High model utility, versatility, efficiency, fine-tuning support Resource intensive, overfitting, high maintenance and updates
Hugging Face transformers [HuggingFaceTransformers] Hugging Face Open source NLP, model fine-tuning, text generation, research and development large model repository, extensive community support, interoperability, versatility No official support, resource intensive
Google Cloud AI-Language (Gemini, Bard, PaLM, etc.) [GoogleCloudNLP] Google API Sentiment analysis, entity recognition, translation, dialogue, text/code/image generation, comprehension Google’s robust and substantial infrastructure, easy integration with Google services, proficiency Pricing based on usage, quality of generated content, inconsistent responses
Amazon Comprehend [satyanarayana2020sentimental] Amazon AWS API Entity recognition, sentiment analysis, classification, topic modeling Integration with other AWS services, scalability, reliability Pricing based on usage, limited customization
IBM Watson NLU [IBMNLP] IBM Watson API Sentiment and emotion analysis, keyword extraction Comprehensive text analysis, customization, scalability, rich documentation and support, Integration with other IBM services Cost for high usage, overhead for simple tasks
Fairseq [satyanarayana2020sentimental] Meta Open source Translation, research and development, language modeling Versatility, research-friendly, integration with PyTorch Limited user community, less user-friendly for beginners, maintenance
Claude [AnthropicClaude] Anthropic API Summarization, search, collaborative writing, coding comprehension Safety and alignment focus, human-like interaction, multi-modality Costly, limited customization
Cohere [Cohere] Cohere Inc. API Chatbots, knowledge assistants, scalability User-friendly, seamless customization, focus on enterprise needs, privacy Costly, small community
TABLE VII: Details on WM Schemes used in Experiments.
WM Details Mechanism Visualization
KGW [kirchenbauer2023watermark] Divides the vocabulary into green and red tokens, then modifies logits by adding a fixed WM to green tokens’ logits to favor the generation of green tokens. [Uncaptioned image]
SIR [liu2024semanticinvariantrobustwatermark] Determines WM logits by converting semantics of all preceding tokens using an auxiliary embedding model, then combines them with the original logits. [Uncaptioned image]
EXP [kuditipudi2023robust] Intervenes the sampling process of each token by mapping it with a pseudo-random number sequence, in which every generated token corresponds with a key. [Uncaptioned image]
SemStamp [hou2023semstamp] Generates new sentences by mapping candidates into embeddings, partitions semantic space into valid & blocked regions, and chooses sentences in valid regions. [Uncaptioned image]
Unigram [zhao2024provable] Modifies the logits to embed the WM by using a fixed grouping strategy, same green and red lists applied consistently for entire vocabulary. [Uncaptioned image]
Adaptive [liu2024adaptive] Adaptively scales up logits of high-entropy tokens based on preceding tokens’ semantics embedding. [Uncaptioned image]
UPV [liu2023unforgeable] Employs two networks for WM generation (modifying logits) and detection (evaluating entire text without a generation key). [Uncaptioned image]